NASA Jet Propulsion Laboratory California Institute of Technology JPL HOME EARTH SOLAR SYSTEM STARS & GALAXIES SCIENCE & TECHNOLOGY JPL Email News RSS Mobile Video
Follow this link to skip to the main content
JPL banner - links to JPL and CalTech
left nav graphic Overview Science Technology The Mission People Spotlights Events Multimedia All Mars
Mars for Kids
Mars for Students
Mars for Educators
Mars for Press
+ Mars Home
+ Rovers Home
image link to mission page
image link to summary page
link to rovers update page
Spirit Archive
Opportunity Archive
Where are they now?
month in review
image link to mission team
image link to launch vehicle
image link to spacecraft
link to mission timeline page
communications to earth
Update: Spirit and Opportunity

M I S S I O N     M A N A G E R S   
Scott Lever, Mission manager Mike Seibert, Mission manager Al Herrera, Mission manager
Scott Lever Mike Seibert Al Herrera

SPIRIT UPDATE:  Spirit Remains Silent at Troy - sols 2621-2627, May 18-24, 2011:

More than 1,300 commands were radiated to Spirit as part of the recovery effort in an attempt to elicit a response from the rover. No communication has been received from Spirit since Sol 2210 (March 22, 2010). The project concluded the Spirit recovery efforts on May 25, 2011. The remaining, pre-sequenced ultra-high frequency (UHF) relay passes scheduled for Spirit on board the Odyssey orbiter will complete on June 8, 2011.

Total odometry is unchanged at 7,730.50 meters (4.80 miles).

Spirit Update Archive


OPPORTUNITY UPDATE:  Opportunity's Heading to a Small Crater Called 'Ulysses' - sols 3786-3792, September 17, 2014-September 23, 2014:

Opportunity is on the west rim of Endeavour Crater heading towards "Marathon Valley," a putative location for abundant clay minerals.

The rover is headed to a near-term target, a small crater named "Ulysses." The rover is moving closer to Ulysses to get a peek inside. On Sol 3787 (Sept. 18, 2014), Opportunity drove a little over 44 feet (13.5 meters) in rocky terrain, requiring the use of Visual Odometry to safely navigate. On Sol 3789 (Sept. 20, 2014), the rover moved closer to the rim of Ulysses, but the drive stopped after 15 feet (4.6 meters) because Visual Odometry was not tracking on the last steps. An evening Alpha Particle X-ray Spectrometer measurement of atmospheric argon was performed on Sol 3790 (Sept. 21, 2014). The rover continued closer to Ulysses on the next sol with a 13-feet (4-meter) bump. High slip prevented the rover from completing the turn for communication at the end of the drive.

Recently, there were more Flash-related events. Two more "amnesia" events occurred on the evenings of Sols 3786 and 3789 (Sept. 17 and Sept. 20, 2014). And two Flash write errors to Bank 7 occurred on Sols 3791 and 3792 (Sept. 22 and Sept. 23, 2014). All these events were benign and did not impact the rover's operation. The project continues to investigate. Otherwise, Opportunity continues in good health.

As of Sol 3792 (Sept. 23, 2014), the solar array energy production was 639 watt-hours with an atmospheric opacity (Tau) of 0.889 and a solar array dust factor of 0.740.

Total odometry is 25.34 (40.77 kilometers).

Opportunity Update Archive

USA.gov
PRIVACY    |     FAQ    |     SITEMAP    |     CREDITS